# Consumption Partial Insurance

### v3

@@ C.2 Sources of tail income shocks
Household income growth
bottom 10% top 90% difference
empirical probabilities $p$-value
@@
Unemployed $t$ (male) 0.082 0.034 0.000
Unemployed $t-1$ (male) 0.084 0.036 0.000
Unemployed $t$ (female) 0.081 0.037 0.000
Unemployed $t-1$ (female) 0.081 0.041 0.000
Main job laid off (male) 0.054 0.039 0.003
Second job laid off (male) 0.153 0.117 0.000
Main job laid off (female) 0.073 0.059 0.023
Second job laid off (female) 0.096 0.095 0.921
Physical disability (male) 0.130 0.103 0.001
Physical disability (female) 0.121 0.120 0.840
Physical disability (both) 0.032 0.025 0.077
Overall health got worse (male) 0.129 0.110 0.036
Overall health got worse (female) 0.121 0.104 0.052
Eviction from home 0.049 0.031 0.000
@@
Notes: The table reports the empirical probability of the various events listed in the rows of the table in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$. Given the biennial nature of the PSID, $\Delta y_{it}$ corresponds to a difference over two calendar years.
C.3 CEX data and imputation in the PSID
The Consumer Expenditure Survey accounts for about $95\%$ of all household expenditures from a highly disaggregated list of consumption goods and services (see Meyer and Sullivan, 2023, for a recent overview). The CEX is run by the Census Bureau and the Bureau of Labor Statistics, and it forms the basis for the calculation of the consumer’s price index.
Sample selection and variables. To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019.333To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID. We select a sample of households that mimics closely our baseline selection in the PSID. Specifically, we select continuously married couples with the male spouse aged 30 to 65. We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the consumption imputation subsequently. Similar to the consumption measure in the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare. As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters. We assign the final observation to the calendar year to which the underlying quarterly data mostly correspond to. Our final sample consists of 31,751 observations.
@@ C.3 CEX data and imputation in the PSID
Figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID (blue solid line) and CEX samples (black dotted line). Although the patterns for the variance and kurtosis are quite similar in the two surveys over time, consumption in the PSID is less volatile relative to the CEX but slightly more leptokurtic. For skewness, there is disagreement both in the trends over time and, importantly, with respect to the sign, with the PSID and CEX samples featuring positive and weakly negative skewness, respectively. Overall, these statistics (in particular the skewness statistics) suggest that consumption across otherwise similarly selected samples in the PSID and CEX exhibits notable differences; this calls into question any conclusions on the pass-through drawn from methods that use higher-order moments of consumption to identify the response to income shocks.
@@
Notes: The figure plots the moments of log consumption in the PSID (blue solid line; left axis) and the CEX (black dotted line for the true series and black dashed-dotted line for the imputed series; right axis).
The imputation involves two distinct steps. In the first step, we estimate food demand in the CEX. Specifically, we let real food expenditure be a log-linear function of the log of nondurable expenditure $c$, relative prices $\mathbf{p}$, and demographics $\mathbf{W}$, namely
@@
where $f_{it}$ is the log of real annual food expenditure for household $i$ in the CEX in year $t$, $\beta$ is the budget elasticity of food demand, which we allow to shift with time and household characteristics $D_{it}$, and $e_{it}$ captures unobserved heterogeneity in food demand. [^5] Prices $\mathbf{p}$ include the relative prices of food, public transportation, utilities, medical services, and childcare. Demographics $\mathbf{W}$ include a quadratic polynomial in age and dummy variables for the number of children, family size, education, region of residence, race, and cohort. $D_{it}$ includes year dummies and dummies for education and the number of children. To address possible endogeneity of and measurement error in total expenditure, we follow an instrumental variables approach using as instruments the mean and standard deviation of household disposable income and the standard deviation of spousal wages by cohort, education, year, the former interacted with time, education, and number of children dummies. [^6] These instruments are slightly different from those used in BPP, namely the cohort-education-year specific averages of male and female hourly wages, interacted with time, education, and kids dummies. The use of their instruments over our sample period often resulted in convergence problems.
In the second step, we invert (C.1) under the assumption that food demand is monotonic in total expenditure. We can then use the inverted equation to predict nondurable expenditure $c$ for a given level of food expenditure, given prices and demographics. Since the CEX and PSID samples are representative of the same underlying population, and since food expenditure, prices, and demographics are available in both surveys, we can use the inverted (C.1) to impute consumption into the PSID. This allows us to obtain an alternative, external measure of consumption in the PSID, and thus of consumption growth.
@@
An explanation for why the imputation inflates $\boldsymbol{\phi_{t}^{(1)}}$. Our empirical findings suggest that estimation of the linear specification using imputed consumption data inflates $\phi_{t}^{(1)}$ relative to estimation using the true PSID series. One reason behind this (though likely not the only one) is neglected correlation between measurement error in income and expenditure.
Suppose that consumption and income measurement errors are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{c},u_{it}^{y})=\sigma_{u_{t}}>0$. The positive correlation may be the result of, e.g., recall bias that affects similarly all monetary amounts. This is something we have not allowed for in the paper, and neither have BPP. While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors). This is what we also do in our baseline estimation of $\phi_{t}^{(1)}$ in the linear specification. However, if the errors are correlated, then this target moment in fact identifies $\phi_{t}^{(1)\text{psid*}}=\phi_{t}^{(1)}+\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. Clearly, correlation between consumption and income errors causes positive bias even in the baseline, i.e., without the use of imputed consumption, as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$. We show below that this bias is multiplicatively larger when one uses imputed consumption.
@@
Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{c})$. We assume for simplicity that measurement error in food expenditure is the same as measurement error in nondurable consumption. This is not crucial for what follows; what is crucial is that measurement error in income varies jointly with measurement error in food expenditure in a similar way as it does with measurement error in nondurable consumption. Because $\kappa_{t}>1$, the consumption imputation magnifies the role of the consumption measurement error. One can show that, with imputed consumption, the moment that is used to identify $\phi_{t}^{(1)}$ in fact identify $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation in this environment inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
C.4 Comparison using the original BPP data
@@
Second, we calculate biennial growth rates of income and consumption and re-estimate the linear model. This is column 2 in table C.4; this estimation is analogous to the linear function estimation in column 1 of table 4, albeit using the original BPP sample. Note that despite the use of biennial data, $\phi^{(1)}$ and $\psi^{(1)}$ still express the pass-through at annual rates (see appendix B). $\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains qualitatively unaffected. With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years. Some higher frequency shocks will be muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period. In other words, the use of biennial data allows us to only measure an ‘aggregate’ or average pass-through within a given biennial period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
Third, we re-estimate the linear model assuming that the target second-order income and consumption moments are time-invariant. This differs from the original BPP exercise where the moments vary with time, but it is otherwise similar to our main exercise in the paper given that we cannot precisely estimate higher-order moments on a year-by-year basis. The results are in columns 3-4 in table C.4. Time-invariance further reduces the transmission parameter of permanent shocks ($\phi^{(1)}$ drops by another 20% to $0.404$), bringing the pass-through closer to the value we estimate in the modern data. Taking averages of moments over time attenuates the largest income-consumption co-movements; consequently, the model can fit the data using, ceteris paribus, smaller transmission parameters. This is similar to Chatterjee et al. (2021) who also find smaller pass-through rates (higher degree of partial insurance) when they impose time-invariance in the BPP target moments.
There are some caveats comparing the exercise in table C.4 and our results with the most recent data in table 4. First, our main exercise uses consumption data internally available in the PSID, while table C.4 uses consumption imputed from the CEX. We showed earlier that the imputation imparts substantial imputation error. Second, while the estimation in table C.4 does not correct the income moments for measurement error (in line with BPP), our main exercise in table C.4 does so (in line with Blundell et al., 2016). Finally, the use of biennial growth rates does not allow the identification of $\theta$, the moving average parameter of the transitory shock, so the transitory shock is $MA(0)$ in our baseline ($\theta=0$), while it is $MA(1)$ in the original BPP exercise.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, using the original BPP data over 1980–1992 (PSID income data, consumption data imputed from the CEX). The data are available through Blundell et al. (2008)’s online replication package. Columns 1-2 allow the target second-order income and consumption moments to vary with calendar time, while columns 3-4 target said moments imposing time-invariance. Columns 1 and 3 estimate the model using annual growth rates of income and consumption, while columns 2 and 4 use biennial rates. In line with BPP, estimation is done via diagonally weighted GMM; asymptotic standard errors are in parentheses.
@@
C.5 Additional results
Tables C.5 and C.6 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM). As such, these tables accompany tables 3 and 4 in the paper, which report results from equally weighted GMM. The results are qualitatively (and most often numerically) unchanged from those reported in the paper.
@@ C.5 Additional results
Table C.5: Estimates of the income process, baseline sample, optimal weighting
@@
Variance $\sigma_{\zeta}^{2}$ 0.027 0.027 0.025
(0.002) (0.002) (0.002)
Skewness $\gamma_{\zeta}$ central -0.004 -0.004
(0.003) (0.003)
standardized -0.895 -0.997
(0.633) (0.760)
Kurtosis $\kappa_{\zeta}$ central 0.037
standardized 61.636
@@
Variance $\sigma_{v}^{2}$ 0.033 0.033 0.034
(0.002) (0.002) (0.002)
Skewness $\gamma_{v}$ central -0.008 -0.008
(0.003) (0.003)
standardized -1.351 -1.324
(0.517) (0.485)
Kurtosis $\kappa_{v}$ central 0.049
standardized 41.579
@@
Table C.6: Estimates of the consumption function, baseline sample, optimal weighting
@@
Table C.7: Estimates of the consumption function, baseline sample, imputed consumption
@@
Notes: The table presents the estimates of the parameters of the consumption function, imposing homogeneous transmission parameters over the lifecycle/in the cross-section and stationarity of taste heterogeneity and consumption measurement error. Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. All columns use consumption imputed from the CEX (see appendix C.3). Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. This table accompanies table 4 in the paper, which shows the baseline results using consumption internally available in the PSID.
@@
Table C.8 presents estimation results for the linear consumption function in subsamples formed by wealth and education. As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification. Columns 1-4 present parameter estimates among low and high wealth households. $\phi^{(1)}$ is consistently smaller among the wealthier, regardless of which set of moments are targeted, indicating the self-insurance role of assets in that group. $\psi^{(1)}$ is statistically indistinguishable from zero in both groups. Columns 5-8 present parameter estimates among households with and without college education. The results do not present a clear pattern across groups. When only second moments of income and consumption are targeted, $\phi^{(1)}$ is larger for households without college education – $0.225$ versus $0.123$. However, when all moments up to fourth-order are included, $\phi^{(1)}$ is almost identical in the two groups.
@@
By wealth By education
Subsample: Low wealth High wealth No college Some college

### v4

@@ C.3 Sources of tail income shocks
Disposable income growth
Bottom 10% Top 90%
mean mean $p$-value $\Delta$mean
@@
% unemployed $t$ (male) 0.082 0.034 0.000
% unemployed $t-1$ (male) 0.084 0.036 0.000
% unemployed $t$ (female) 0.081 0.037 0.000
% unemployed $t-1$ (female) 0.081 0.041 0.000
% main job laid off (male) 0.054 0.039 0.003
% second job laid off (male) 0.153 0.117 0.000
% main job laid off (female) 0.073 0.059 0.023
% second job laid off (female) 0.096 0.095 0.921
% physical disability (male) 0.130 0.103 0.001
% physical disability (female) 0.121 0.120 0.840
% physical disability (both) 0.032 0.025 0.077
% overall health got worse (male) 0.129 0.110 0.036
% overall health got worse (female) 0.121 0.104 0.052
% eviction from home 0.049 0.031 0.000
@@
Notes: The table reports the sample mean (mostly empirical probabilities) of the various events listed in the rows of the table in the bottom 10% and top 90% of the distribution of residual income growth $\Delta y_{it}$. Given the biennial nature of the PSID, $\Delta y_{it}$ corresponds to a difference over two calendar years.
C.4 CEX data and imputation in the PSID
The Consumer Expenditure Survey (CEX) accounts for about $95\%$ of all household expenditures from a highly disaggregated list of consumption goods and services (see Meyer and Sullivan, 2023, for a recent overview). The CEX is run by the Census Bureau and the Bureau of Labor Statistics and forms the basis for the calculation of the consumer’s price index.
Sample selection and variables. To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019. We select a sample of households that mimics closely our baseline selection in the PSID. [^3] To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID. Specifically, we select continuously married couples with the male spouse aged 30 to 65. We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the imputation subsequently. Similar to the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare. As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters. We assign the final observation to the calendar year to which the underlying quarterly data mostly correspond to. Our final sample consists of 31,751 observations.
@@ C.4 CEX data and imputation in the PSID
Figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID (blue solid line) and CEX samples (black dotted line). The patterns for the variance and kurtosis are quite similar in the two surveys over time, although consumption in the PSID is less volatile but slightly more leptokurtic. For skewness, there is disagreement both in the trends over time and the sign. The PSID features positive while the CEX weakly negative skewness. Overall, these statistics (in particular skewness) suggest that consumption across otherwise similarly selected samples in the PSID and CEX exhibits some notable differences.
@@
Notes: The figure plots the moments of log consumption in the PSID (blue solid line; left), the CEX (black dotted line; right), and imputed from the CEX into the PSID (black dashed-dotted line; right axis).
The imputation involves two distinct steps. In the first step, we estimate food demand in the CEX. Specifically, we let real food expenditure be a log-linear function of the log of nondurable expenditure $c_{it}$, relative prices $\mathbf{p}_{t}$, and demographics $\mathbf{W}_{it}$, namely
@@
where $f_{it}$ is the log of real annual food expenditure for household $i$ in the CEX in year $t$, $\beta$ is the budget elasticity of food demand, which we allow to shift with time and household characteristics $D_{it}$, and $e_{it}$ captures unobserved heterogeneity in food demand. [^5] Prices $\mathbf{p}_{t}$ include the relative prices of food, public transportation, utilities, medical services, and childcare. Demographics $\mathbf{W}_{it}$ include a quadratic polynomial in age and dummy variables for the number of children, family size, education, region of residence, race, and cohort. $D_{it}$ includes year dummies and dummies for education and the number of children. To address possible endogeneity of nondurable expenditure, we follow an instrumental variables approach using as instruments the mean and standard deviation of household disposable income and the standard deviation of spousal wages by cohort, education, year, the former interacted with time, education, and number of children dummies. [^6] These instruments are different from those used in BPP, namely the cohort-education-year specific averages of male and female hourly wages, interacted with time, education, and kids dummies. The use of their instruments over our sample period often resulted in convergence problems.
In the second step, we invert (C.1) under the assumption that food demand is monotonic in nondurable expenditure. We can then use the inverted equation to predict $c_{it}$ for a given level of food expenditure, given prices and demographics. Since the CEX and PSID samples are representative of the same underlying population, and since food expenditure, prices, and demographics are available in both surveys, we can use the inverted (C.1) to impute consumption into the PSID. This allows us to obtain an alternative, external measure of consumption in the PSID, and thus of consumption growth.
@@
An explanation for why the imputation inflates $\boldsymbol{\phi_{t}^{(1)}}$. Our empirical findings suggest that estimation of the linear specification using imputed consumption data inflates $\phi_{t}^{(1)}$ relative to estimation using the true PSID series. One reason behind this is neglected correlation between measurement error in income and expenditure.
Suppose that consumption and income measurement errors are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{c},u_{it}^{y})=\sigma_{u_{t}}>0$. The positive correlation may be the result of, e.g., recall bias that affects similarly all monetary amounts. This is something we have not allowed for in the paper, and neither have BPP. While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors). This is what we also do in our baseline estimation of $\phi_{t}^{(1)}$ in the linear specification. However, if the errors are correlated, then this target moment in fact identifies $\phi_{t}^{(1)\text{psid*}}=\phi_{t}^{(1)}+\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. Clearly, correlation between consumption and income errors causes positive bias as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$. This is true in the baseline, i.e., without the use of imputed consumption, but the bias becomes multiplicatively larger with imputed consumption, as we show below.
@@
Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{f})$. Suppose that food consumption and income error are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{f},u_{it}^{y})=\sigma_{u_{t}}>0$. Because $\kappa_{t}>1$, the consumption imputation magnifies the role of the consumption measurement error. One can show that, with imputed consumption, the moment used to identify $\phi_{t}^{(1)}$ in fact identifies $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
C.5 Comparison using the original BPP data
@@
Second, we calculate biennial growth rates of income and consumption and re-estimate the linear model. This is column 2 in table C.5; this estimation is analogous to the linear function estimation in column 1 of table 3, albeit using the original BPP sample. Note that despite the use of biennial data, $\phi^{(1)}$ and $\psi^{(1)}$ still express the pass-through at annual rates (see appendix B). $\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains qualitatively unaffected. With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years. Some higher frequency shocks are muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period. In other words, the use of biennial data allows us to only measure an ‘aggregate’ or average pass-through within a given biennial period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
Third, we re-estimate the linear model assuming that the target second-order income and consumption moments are time-invariant. This differs from the original BPP exercise where the moments vary with time, but it is otherwise similar to our main exercise in the paper given that we cannot precisely estimate higher-order moments on a year-by-year basis. The results are in columns 3-4 in table C.5. Time-invariance further reduces the transmission parameter of permanent shocks ($\phi^{(1)}$ drops to $0.404$), bringing the pass-through closer to the value we estimate in the modern data. Taking averages of moments over time attenuates the largest income-consumption co-movements; consequently, the model can fit the data using, ceteris paribus, smaller transmission parameters. This is similar to Chatterjee et al. (2021) who also find smaller pass-through rates (higher degree of partial insurance) when they impose time-invariance in BPP’s target moments.
There are some caveats comparing the exercise in table C.5 and our results with the most recent data in table 3. First, our main exercise uses consumption data internally available in the PSID, while table C.5 uses consumption imputed from the CEX. We showed earlier that the imputation imparts substantial imputation error. Second, while the estimation in table C.5 does not correct the income moments for measurement error (in line with BPP), our main exercise in table 3 does so (in line with Blundell et al., 2016). Finally, the use of biennial growth rates does not allow the identification of $\theta$, the moving average parameter of the transitory shock, so the transitory shock is $MA(0)$ in our baseline ($\theta=0$), while it is $MA(1)$ in the original BPP exercise.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, using the original BPP data over 1980–1992 (PSID income data, consumption data imputed from the CEX). The data are available through Blundell et al. (2008)’s online replication package. Columns 1-2 allow the target second-order income and consumption moments to vary with calendar time, while columns 3-4 target said moments imposing time-invariance. Columns 1 and 3 estimate the model using annual growth rates of income and consumption, while columns 2 and 4 use biennial rates. In line with BPP, estimation is done via diagonally weighted GMM; GMM standard errors are in parentheses. The very small unimportant numerical differences between column 1 herein and the original BPP are likely due to different software and optimization algorithms (we use Matlab and fmincon).
@@
C.6 Additional results
Tables C.6 and C.7 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM). As such, these tables accompany tables 2 and 3 in the paper, which report results from equally weighted GMM. The results are qualitatively (and most often numerically) similar to those reported in the paper.
@@ C.6 Additional results
Table C.9 presents the estimates of the parameters of the income process by age and the household’s position in the income distribution. Skewness of all shocks becomes more negative, both economically and statistically, with age. Skewness of permanent shocks is weakly positive in the bottom income tertile but turns very negative in the top tertile. The opposite is true for transitory shocks; skewness is negative in the bottom income tertile but strongly positive in the top.
Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution. Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted. This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^7] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets. Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, possess more insurance than the poor.
Table C.6: Estimates of the income process, optimal weighting
@@
Variance $\sigma_{\zeta}^{2}$ 0.027 (0.002) 0.027 (0.002) 0.025 (0.002)
Skewness $\gamma_{\zeta}$ central -0.004 (0.003) -0.004 (0.003)
standardized -0.895 (0.633) -0.997 (0.760)
Kurtosis $\kappa_{\zeta}$ central 0.037 (0.007)
standardized 61.636 (14.018)
@@
Variance $\sigma_{v}^{2}$ 0.033 (0.002) 0.033 (0.002) 0.034 (0.002)
Skewness $\gamma_{v}$ central -0.008 (0.003) -0.008 (0.003)
standardized -1.351 (0.517) -1.324 (0.485)
Kurtosis $\kappa_{v}$ central 0.049 (0.005)
standardized 41.579 (7.052)
@@
Table C.7: Estimates of the consumption function, optimal weighting
@@
Table C.8: Estimates of the consumption function, imputed consumption
@@
Notes: The table presents the estimates of the parameters of the consumption function, imposing homogeneous transmission parameters over the lifecycle/in the cross-section and stationarity of taste heterogeneity and consumption measurement error. Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. All columns use consumption imputed from the CEX (see appendix C.4). Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. This table accompanies table 3 in the paper, which shows the baseline results using consumption internally available in the PSID (column 2 in table 3 is the same as column 1 in the present table).
@@
Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education. As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification. The odd numbered columns present parameter estimates when only second-order moments are targeted. $\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across the groups. This is similar to what we found in the quadratic specification. However, as soon as we target up to fourth-order moments in the linear specification, shown in the even numbered columns, this monotonic pattern is lost. Transitory shocks transmit strongly in group 1, and with a negative coefficient.
Table C.9: Estimates of the income process, by age and position in income distribution
Age Position in income distribution $y_{it-1}$
30–40 41–50 51–65 Bottom tertile Top tertile
Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th
Income data: PSID PSID PSID PSID PSID
(1) (2) (3) (4) (5)
Panel A. Permanent shocks
Variance $\sigma_{\zeta}^{2}$ 0.028 0.026 0.034 0.037 0.032
(0.004) (0.004) (0.005) (0.004) (0.003)
Skewness $\gamma_{\zeta}$ central -0.003 0.004 -0.015 0.010 -0.028
(0.006) (0.005) (0.006) (0.005) (0.004)
standardized -0.653 0.926 -2.386 1.344 -4.851
(1.291) (1.114) (1.167) (0.783) (0.987)
Kurtosis $\kappa_{\zeta}$ central 0.043 0.030 0.058 0.052 0.039
(0.014) (0.013) (0.016) (0.013) (0.008)
standardized 56.275 45.537 50.945 37.627 36.974
(23.787) (22.709) (21.605) (13.058) (10.586)
Panel B. Transitory shocks
Variance $\sigma_{v}^{2}$ 0.033 0.027 0.036 0.020 0.020
(0.005) (0.004) (0.005) (0.004) (0.003)
Skewness $\gamma_{v}$ central -0.008 -0.005 -0.010 -0.024 0.021
(0.006) (0.005) (0.005) (0.005) (0.004)
standardized -1.427 -1.178 -1.453 -8.246 7.333
(1.057) (1.088) (0.782) (3.074) (2.235)
Kurtosis $\kappa_{v}$ central 0.056 0.040 0.050 0.014 0.011
(0.013) (0.007) (0.008) (0.008) (0.006)
standardized 52.863 53.730 38.832 33.974 27.286
(19.670) (16.619) (12.566) (24.531) (16.935)
Notes: The table presents the estimates of the parameters of the income process, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse, and, separately, by the household’s position in the income distribution (bottom and top tertiles of the distribution of lag income $y_{it-1}$). We target all income moments up to fourth-order; results when targeting only lower order moments are available upon request. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies tables 4 and 5 in the paper, as well as the following tables C.10 and C.11.
Table C.10: Estimates of the linear consumption function, by age
Consumption fn.: Linear
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
Consumption data: PSID PSID PSID
(1) (2) (3)
$\phi^{(1)}$ 30-40 0.154 (0.045) 0.164 (0.051) 0.089 (0.058)
41-50 0.126 (0.051) 0.146 (0.055) 0.079 (0.089)
51-65 0.180 (0.049) 0.175 (0.051) 0.136 (0.061)
$\psi^{(1)}$ 30-40 0.094 (0.079) 0.083 (0.073) 0.095 (0.107)
41-50 0.010 (0.079) -0.016 (0.072) -0.075 (0.061)
51-65 -0.046 (0.070) -0.044 (0.066) 0.037 (0.101)
$\sigma^{2}_{\xi}$ 30-40 0.014 (0.002) 0.014 (0.002) 0.016 (0.002)
41-50 0.022 (0.002) 0.022 (0.002) 0.022 (0.002)
51-65 0.021 (0.002) 0.021 (0.002) 0.023 (0.002)
$\gamma_{\xi}$ 30-40 -0.001 (0.001) -0.001 (0.001)
41-50 0.006 (0.001) 0.006 (0.002)
51-65 -0.006 (0.002) -0.006 (0.002)
$\kappa_{\xi}$ 30-40 0.004 (0.002)
41-50 0.007 (0.002)
51-65 0.006 (0.002)
$\sigma^{2}_{u_{c}}$ 30-40 0.046 (0.003) 0.046 (0.003) 0.046 (0.003)
41-50 0.040 (0.002) 0.040 (0.002) 0.040 (0.002)
51-65 0.043 (0.002) 0.043 (0.002) 0.043 (0.002)
$\gamma_{u_{c}}$ 30-40 0.000 (0.002) 0.000 (0.002)
41-50 0.002 (0.002) 0.002 (0.002)
51-65 0.004 (0.002) 0.004 (0.002)
$\kappa_{u_{c}}$ 30-40 0.014 (0.003)
41-50 0.012 (0.002)
51-65 0.012 (0.002)
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse. The underlying income moments also vary over these age brackets (table C.9). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. This table accompanies table 4 in the paper, which shows results from the quadratic specification.
Table C.11: Estimates of the linear consumption function, over position in income distribution
Consumption fn.: Linear
Lag income $y_{it-1}$: Bottom tertile Top tertile
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd 2nd, 3rd 2nd, 3rd, 4th
Consumption data: PSID PSID PSID PSID PSID PSID
(1) (2) (3) (4) (5) (6)
$\phi^{(1)}$ 0.103 0.083 0.063 0.102 0.060 0.036
(0.031) (0.032) (0.035) (0.036) (0.034) (0.041)
$\psi^{(1)}$0.045 0.013 0.025 -0.004 0.101 0.071
(0.086) (0.038) (0.054) (0.093) (0.048) (0.073)
$\sigma^{2}_{\xi}$ 0.014 0.014 0.014 0.025 0.025 0.026
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
$\gamma_{\xi}$0.001 -0.001 0.002 0.002
(0.001) (0.001) (0.001) (0.001)
$\kappa_{\xi}$ 0.003 0.009
(0.002) (0.002)
$\sigma^{2}_{u_{c}}$ 0.044 0.044 0.044 0.041 0.041 0.041
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
$\gamma_{u_{c}}$ 0.003 0.003 0.003 0.003
(0.002) (0.002) (0.002) (0.002)
$\kappa_{u_{c}}$ 0.011 0.010
(0.002) (0.002)
Notes: The table presents the estimates of the parameters of the linear consumption function in the bottom and top tertiles of the distribution of lag income $y_{it-1}$. The underlying income moments also vary over the income distribution (table C.9). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies table 5 in the paper, which shows results from the quadratic specification.
@@
Low wealth High wealth
Subsample: No college Some college No college Some college